# On Inverse Problems, Parameter Estimation, and Domain Generalization

### v1

@@ A.3 Minimal Sufficient Statistic
Theorem 15 (sufficient statistic (Cover and Thomas, 2006))
@@
A.4 Data Processing Inequality
@@
$H(\theta|X)=H(\theta|X,T(X))\leq H(\theta|T(X)),$ (60)
@@
A.5 Fisher-Neyman Factorization Theorem
@@
$\frac{\partial}{\partial\theta}\log(p_{\theta}(x))=\frac{\partial}{\partial\theta}\log(q_{\theta}(T(x))).$ (62)
Therefore, $J_{X}(\theta)=J_{T(X)}(\theta)$. In other words, $T$ is sufficient if and only if the densities can be written as products of two factors, one factor that depends on the outcome through $T$ only, and a second factor that is independent of the unknown measure (Halmos and Savage, 1949). Another way of phrasing this result is to say that $T$ is sufficient if and only if the likelihood ratio of every pair of measures depends on the outcome through $T$ only.
@@
When sufficient statistics exist it has been shown that they will be solutions of the equations of Maximum Likelihood (Fisher, 1925).
A.6 The Rao-Blackwell Theorem
Given an unbiased estimator $f(X)$ and a sufficient statistic $T(X)$, a Rao-Blackwell estimator is formed by conditioning $f(X)$ on the sufficient statistic $T(X)$, i.e., $Ef(X)|T(X)$. This estimator does not depend on $\theta$, and will have a smaller variance or be as efficient as any other unbiased estimator (Blackwell, 1947). That said, note that not all sufficient statistics lead to efficient estimators that achieve the CRB. The efficiency of an estimator also depends on other factors, such as the specific form of the estimator and the underlying distribution of the data.
A.7 Estimation uncertainty bounds
Theorem 16 (Estimation error and differential entropy (Cover and Thomas, 2006))
@@ A.7 Estimation uncertainty bounds
[…]
Appendix B Sparse Recovery Method and Recovery-Error Bound
@@
$x[k]=\sum_{m}{c_{m}\delta[k-k_{m}]},\qquad k\in\mathbb{Z},\qquad c_{m}\in\mathbb{R}$ (64)
where $\delta[k]$ denotes the Kronecker delta function (see (Ricker, 1940)), and $\sum_{m}|c_{m}|<\infty$. $K=\{k_{m}\}$ is the set of discrete delays corresponding to the spikes locations. In a noisy environment we consider a discrete signal of the form
$y[k]=\sum_{n}{x[n]g_{\sigma,n}[k-n]}+w[k],\qquad n\in\mathbb{Z}$ (65)
where $w[k]$ is additive noise, with bounded energy $\|\mathbf{w}\|_{1}=\|\mathbf{y}-G\mathbf{x}\|_{1}\leq\delta,$ where $\|\mathbf{x}\|_{1}:=\sum_{k}|x[k]|$ and $G$ as an operator matrix such that $(G)_{k,n}=g_{\sigma,n}[k-n]$, representing a set of admissible kernels with parameters $\beta,\varepsilon$ as defined in (Pereg and Cohen, 2017, Definition 2.1 ), and a scaling parameter $\sigma$. We denote $F_{\mathrm{s}}$ as the sampling rate. The objective is to estimate the true support $K=\{k_{m}\}$ and the spikes’ amplitudes $\{c_{m}\}$ from the observed signal (e.g., seismic trace) $y[k]$. In (Pereg and Cohen, 2017) it is shown that the solution $\hat{\mathbf{x}}$ of
${\begin{array}[]{lll}&\min&\|\mathbf{\hat{x}}\|_{1}\\ &{\text{subject to}}&\|\mathbf{y}-G\mathbf{\hat{x}}\|_{1}\leq\delta,\end{array}}$ (68)
@@
$\|\hat{\mathbf{x}}-\mathbf{x}\|_{1}\leq\frac{4\rho}{\beta\gamma_{0}}\delta.$ (69)
@@
By assumption eq. (68) defines a MAP solution where it is assumed that both the solution $\mathbf{x}$ and the noise $\mathbf{w}$ are i.i.d with Laplace zero-mean distribution. Following the same proof outline in (Pereg and Cohen, 2017) it is possible to show (Pereg et al., 2019) that the solution $\hat{\mathbf{x}}$ of
${\begin{array}[]{lll}&\min&\|\mathbf{\hat{x}}\|_{1}\\ &{\text{subject to}}&\|\mathbf{y}-G\mathbf{\hat{x}}\|_{2}\leq S_{\mathrm{w}},\end{array}}$ (72)
@@
$\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}\leq\frac{64N\rho^{2}}{\beta^{2}\gamma^{2}_{0}}S_{\mathrm{w}}.$ (73)

### v2

@@ A.4 Minimal Sufficient Statistic
Theorem A.4 (sufficient statistic [ThomasCover:2006]).
@@
A.5 Data Processing Inequality
@@
(66) $H(\theta|X)=H(\theta|X,T(X))\leq H(\theta|T(X)),$
@@
A.6 Fisher-Neyman Factorization Theorem
@@
(68) $\frac{\partial}{\partial\theta}\log(p_{\theta}(x))=\frac{\partial}{\partial\theta}\log(q_{\theta}(T(x))).$
Therefore, $J_{X}(\theta)=J_{T(X)}(\theta)$. In other words, $T$ is sufficient if and only if the densities can be written as products of two factors, one factor that depends on the outcome through $T$ only, and a second factor that is independent of the unknown measure [Halmos:1949]. Another way of phrasing this result is to say that $T$ is sufficient if and only if the likelihood ratio of every pair of measures depends on the outcome through $T$ only.
@@
When sufficient statistics exist it has been shown that they will be solutions of the equations of Maximum Likelihood [Fisher:1925].
A.7 The Rao-Blackwell Theorem
Given an unbiased estimator $f(X)$ and a sufficient statistic $T(X)$, a Rao-Blackwell estimator is formed by conditioning $f(X)$ on the sufficient statistic $T(X)$, i.e., $Ef(X)|T(X)$. This estimator does not depend on $\theta$, and will have a smaller variance or be as efficient as any other unbiased estimator [Blackwell:1947]. That said, note that not all sufficient statistics lead to efficient estimators that achieve the CRB. The efficiency of an estimator also depends on other factors, such as the specific form of the estimator and the underlying distribution of the data.
A.8 Estimation uncertainty bounds
Theorem A.5 (Estimation error and differential entropy [ThomasCover:2006]).
@@ A.8 Estimation uncertainty bounds
$h(X)$ is the differential entropy of a continuous source $x$ with density $p(x)$ defined as $h(X)\triangleq-\int p(x)\log p(x)dx$. For any random variable with variance $\beta$, $h(X)\leq\frac{1}{2}\mathrm{ln}(2\pi e\beta^{2})$.
Appendix B Proof of Theorems 3.2
Proof B.1.
The model $\mathcal{F}$ is trained by empirical risk minimization
(70) $\mathcal{R}_{m}(\mathcal{F})=\frac{1}{m}\sum_{k=1}^{m}\ell(\mathcal{F}(\mathbf{y}_{k}),\mathbf{x}_{k}).$
where $m$ is the sample size. For $m\rightarrow\infty$, $\mathcal{R}_{m}(\mathcal{F})\rightarrow\mathcal{R}(\mathcal{F})$,
(71) $\small\mathcal{R}(\mathcal{F})=E_{\theta,\mathbf{x},\mathbf{y}}\big[\ell(\mathcal{F}(\mathbf{y}),\mathbf{x})\big]=\frac{1}{M}\sum_{i=1}^{M}E_{\mathbf{x},\mathbf{y}|\theta_{i}}\big[\ell(\mathcal{F}(\mathbf{y}),\mathbf{x})\big]=E_{\mathbf{y}}\Bigg[\frac{1}{M}\sum_{i=1}^{M}E_{\mathbf{x}|\mathbf{y},\theta_{i}}\big[\ell(\mathcal{F}(\mathbf{y}),\mathbf{x})\big]\Bigg].$}
Fix an arbitrary $\mathbf{y}$ in the joint support, since $\mathcal{F}$ operates pointwise on $\mathbf{y}$, minimizing the global risk is equivalent to minimizing the inner term for each $\mathbf{y}$, thus
(72) $\hat{\mathbf{x}}=\mathcal{F}(\mathbf{y})=\arg\min_{\hat{\mathbf{x}}}\frac{1}{M}\sum_{i=1}^{M}E_{\mathbf{x}_{i}\sim\mathbf{x}|\mathbf{y},\theta_{i}}\big[\ell(\hat{\mathbf{x}},\mathbf{x}_{i})\big],$
The theorem formalizes that domain randomization with uniform weighting induces a conditional Bayes estimator under the mixture posterior across domains. If $p(\mathbf{x}|\mathbf{y},\theta_{i})$ collapses to a Dirac delta at $\mathbf{x}_{i}$ for each domain $i$, then $E_{\mathbf{x}_{i}\sim\mathbf{x}|\mathbf{y},\theta_{i}}\big[\ell(\hat{\mathbf{x}},\mathbf{x}_{i})\big]=\ell(\hat{\mathbf{x}},\mathbf{x}_{i})$ and we have
(73) $\hat{\mathbf{x}}=\mathcal{F}(\mathbf{y})=\arg\min_{\hat{\mathbf{x}}}\frac{1}{M}\sum_{i=1}^{M}{\ell(\hat{\mathbf{x}},\mathbf{x}_{i})}.$
Appendix C Sparse Recovery Method and Recovery-Error Bound
@@
(74) $x[k]=\sum_{m}{c_{m}\delta[k-k_{m}]},\qquad k\in\mathbb{Z},\qquad c_{m}\in\mathbb{R}$
where $\delta[k]$ denotes the Kronecker delta function (see [Ricker:1940]), and $\sum_{m}|c_{m}|<\infty$. $K=\{k_{m}\}$ is the set of discrete delays corresponding to the spikes locations. In a noisy environment we consider a discrete signal of the form
(75) $y[k]=\sum_{n}{x[n]g_{\sigma,n}[k-n]}+w[k],\qquad n\in\mathbb{Z}$
where $w[k]$ is additive noise, with bounded energy $\|\mathbf{w}\|_{1}=\|\mathbf{y}-G\mathbf{x}\|_{1}\leq\delta,$ where $\|\mathbf{x}\|_{1}:=\sum_{k}|x[k]|$ and $G$ as an operator matrix such that $(G)_{k,n}=g_{\sigma,n}[k-n]$, representing a set of admissible kernels with parameters $\beta,\varepsilon$ as defined in [Pereg:2017, Definition 2.1 ], and a scaling parameter $\sigma$. We denote $F_{\mathrm{s}}$ as the sampling rate. The objective is to estimate the true support $K=\{k_{m}\}$ and the spikes’ amplitudes $\{c_{m}\}$ from the observed signal (e.g., seismic trace) $y[k]$. In [Pereg:2017] it is shown that the solution $\hat{\mathbf{x}}$ of
(78) ${\begin{array}[]{lll}&\min&\|\mathbf{\hat{x}}\|_{1}\\ &{\text{subject to}}&\|\mathbf{y}-G\mathbf{\hat{x}}\|_{1}\leq\delta,\end{array}}$
@@
(79) $\|\hat{\mathbf{x}}-\mathbf{x}\|_{1}\leq\frac{4\rho}{\beta\gamma_{0}}\delta.$
@@
By assumption eq. (78) defines a MAP solution where it is assumed that both the solution $\mathbf{x}$ and the noise $\mathbf{w}$ are i.i.d with Laplace zero-mean distribution. Following the same proof outline in [Pereg:2017] it is possible to show [Pereg:2019] that the solution $\hat{\mathbf{x}}$ of
(82) ${\begin{array}[]{lll}&\min&\|\mathbf{\hat{x}}\|_{1}\\ &{\text{subject to}}&\|\mathbf{y}-G\mathbf{\hat{x}}\|_{2}\leq S_{\mathrm{w}},\end{array}}$
@@
(83) $\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}\leq\frac{64N\rho^{2}}{\beta^{2}\gamma^{2}_{0}}S_{\mathrm{w}}.$